# AlignMMBench: Evaluating Chinese Multimodal Alignment in Large Vision-Language Models

### v2

@@ 3.1 Dataset composition
  • Incoherence. When a model provides incorrect or irrelevant responses, users usually highlight these errors and expect the model to correct its replies. This task assesses the model’s self-correcting abilities when users indicate its mistakes.
3.2 Dataset construction
As illustrated in Figure 2, the construction of AlignMMBench involves three primary steps: image collection, query generation, and answer annotation. This section provides a concise overview of these steps.
Image collection.
To establish a diverse and high-quality benchmark, we collect images through two primary methods: web image crawlers and real-world user requests. Each image underwent rigorous manual screening to ensure suitability for public use and to avoid inclusion in known SFT datasets or prior benchmarks.
Query generation.
For images obtained via the web crawler, we craft a seed question that correlates with the image and aligns with its designated category. Given the complexity of constructing a dialogue history based on an image, we use real-world user requests exclusively for building the "Dialogue Context" category. Considering the variability of problems with the same user intention, we employ ChatGLM [12] to rephrase the seed questions without altering their original purpose. This method was applied specifically to single-turn questions.
Answer annotation.
We adopt a two-phase annotation process. First, human annotators craft responses based on the corresponding images. Second, researchers review these preliminary annotations, filter out low-quality entries, and request re-annotations where necessary.
4 CritiqueVLM
Figure 3: Critical examples from CritiqueVLM.
In this section, we develop a LLM-based evaluator, referred to as CritiqueVLM , to automatically assess responses based on the questions and detailed reference answers. In addition to assigning a score between 1 and 10, CritiqueVLM also provides a chain-of-thought explanation, shown in Figure 3. CritiqueVLM is fine-tuned from ChatGLM3-6B, an open-source model in the ChatGLM3 series [12, 42].
4.1 Training data construction
Prompt design.
Following  [43], we devise a prompt system consisting of general and task-specific prompts. The general prompt outlines the scoring range, criteria, and output format. Task-specific prompts are varied by different categories, containing several hints to guide the critic model in accurately checking answers. Besides, we use in-context examples to mention that the critic model cannot access an image, encouraging it to infer the image’s content. The detailed prompts are presented in our supplementary materials.
Scoring data collection.
We collect the training corpus based on CogVLM [35] and QwenVL-Chat [4]. Given the challenging nature of the test cases in our AlignMMBench even for human annotators, we utilize responses from a "reference model" that rephrased reference answers using GLM-4 [12, 42]. This approach mitigates bias from predominantly low scores. Human annotators were then tasked with scoring each response from 1 to 10 and providing reasonable explanations, guided by the same prompts used by CritiqueVLM , but with access to images, questions, and reference answers. To prevent bias from the models’ capabilities, we anonymize the model names and randomly shuffled the data.
Dataset partition.
We allocate 90% of the data from each category to the training dataset and reserve the remaining 10% as an internal testing dataset. To assess the model’s agreement with humans when analyzing responses from a novel VLM, we create an external testing dataset using responses generated by GPT-4 [2] and annotated scores provided by human experts.
4.2 Supervised fine-tuning
We select ChatGLM3-6B [12, 42] as the base model and fine-tune it on the training dataset. The DeepSpeed [31] library was employed as our training framework, utilizing the FusedEmaAdam optimizer with a weight decay of 0.05. The learning rate is initially set to $10^{-5}$ with a 1% warmup ratio. The maximum source sequence length is set to 4096, and the target sequence length to 1024. Training is conducted with a global batch size of 128 on 32 A800 GPUs. After 1,000 iterations, the training loss decreased from 3.8 to 0.3.
5 Experiments
Table 2: Mean absolute error ($e$), Pearson ($r$), Spearman ($\rho$), Kendall ($\tau$) correlations, Fuzzy division scores ($f$) and Strict division scores ($s$) are calculated on the internal and external test sets. The best performance among the methods is bold.
Testcase Internal External
Metric $e$ $r$ $\rho$ $\tau$ $f$ $s$ $e$ $r$ $\rho$ $\tau$ $f$ $s$
5.1 Baselines
We select two models, ChatGPT (gpt-3.5-turbo) and GPT-4 (gpt-4-1106-preview) [2], as baselines of CritiqueVLM , which are widely used in other benchmarks. Additionally, given AlignMMBench is a Chinese benchmark, ChatGLM3-6B [12, 42] is also chosen as a baseline.
5.2 Main results
In our experiment, we employ six statistical metrics to assess the agreement between human annotated scores and model generated scores. Common four statistical metrics include the mean absolute error ($e$), Pearson ($r$), Spearman ($\rho$), and Kendall correlation coefficient ($\tau$). To mitigate the bias from annotators preferences, we map scores to predefined ranges and calculate accuracy as an evaluation metric. Based on scoring criteria, integers from 1 to 10 are divided into two sets of ranges:
  • Fuzzy division. This division includes 4 ranges: Unfollow ($[1,2]$), Bad ($[3,5]$), Good ($[6,8]$), and Excellent ($[9,10]$). We denote the proportion of the score from annotation and model in the same ranges as $f$.
  • Strict division. This division includes 7 ranges, each corresponding to specific ranges defined in the scoring criteria: $[1,1],[2,2],[3,3],[4,5],[6,6],[7,8],[9,10]$. We denote the proportion of the score from annotation and model in the same ranges as $s$.
6 Leaderboard
We benchmark a range of popular VLMs, including GPT4o [29], CogVLM2 [35], InternVL-Chat-V1.5 [10, 9], MiniCPM-V2.5 [38, 16, 36, 39], Qwen-VL-Chat [4], XComposer2V [11], DeepSeek-VL [24], Monkey-Chat [19], Yi-VL [3], and Phi-3-Vision [1]. Results are shown in Table 3.
6.1 Analysis on overall result
The average scores of VLMs in Table 3 range from 3.9 to 6.6. According to our scoring criteria, we observe that most VLMs can understand question requirements and generate responses related to the images. Scores below 5.0 indicate numerous errors in the VLMs’ responses. Besides, GPT4o [29] achieves the best performance across most tasks, and CogVLM2 [35] secures the second-best performance. Overall, these scores disparity between GPT4o and other models suggests that GPT4o’s capabilities significantly surpass those of the other models.
6.2 Analysis on category-level performance
Single-turn scenarios.
Focusing on VLMs with average scores above 5.0, we observe that they perform well in tasks such as description, OCR, comparison, and writing. These tasks require VLMs to understand images but do not necessitate complex reasoning or computation. However, they do not perform well in tasks involving Reasoning and Problem-solving. For instance, GPT4o [29] scores only 4.13 in reasoning and 5.85 in problems, indicating frequent response errors. These observations suggest that while current top-tier VLMs can comprehend images and integrate information from images and texts, they struggle with test cases that demand complex reasoning and computation.
Multi-turn scenarios.
In the coherence task, InternVL-Chat-V1.5 [10, 9] demonstrates the best performance among all models. Most VLMs achieve scores above 6.0, indicating their ability to follow instructions from previous interactions and utilize information from the dialogue context effectively. However, all VLMs exhibit suboptimal performance in the incoherence task. This suggests that these models struggle to detect previous errors within the dialogue context and to make accurate corrections as guided by users.
Table 3: Evaluation results on AlignMMBench. Scores in each category are the average scores of test cases. For each column, the highest score is bold, while the second highest score is underlined. The "Ref." column indicates the relative ranking of these models on the https://rank.opencompass.org.cn/leaderboard-multimodal, dominated by primarily English benchmarks.
Models Ref. Avg Perception & Understanding. Reasoning & Analysis Context
Des. Rec. Cou. OCR. Mem. Kno. Rea. Cha. Pro. Com. Wri. Coh. Inc.
GPT-4o [29] 1 6.57 7.69 6.56 6.18 7.4 7.02 6.58 4.13 6.94 5.85 7.09 7.81 6.63 5.5
CogVLM2 [35] 5 5.86 7.19 6.11 5.81 7.03 5.83 5.82 3.57 5.79 4.44 5.65 7.38 6.38 5.22
InternVL-V1.5 [10] 2 5.73 7.21 6.04 5.52 6.64 4.96 5.06 3.34 6.02 4.22 6.41 7.28 6.70 5.07
MiniCPM-V2.5 [16] 3 5.43 7.15 5.32 5.38 6.24 4.55 5.32 3.27 4.78 3.66 6.06 7.34 6.38 5.11
Qwen-VL-Chat [4] 8 5.13 6.43 5.86 5.43 4.77 5.22 5.65 2.90 4.15 3.03 5.52 7.24 6.04 4.46
XComposer2V [11] 4 4.98 6.29 4.71 5.23 4.98 4.69 5.09 3.14 4.43 3.37 4.81 7.17 6.19 4.67
DeepSeek-VL [24] 7 4.75 6.52 5.52 5.14 3.95 3.92 4.18 2.46 3.98 2.61 5.47 7.16 6.17 4.63
Monkey-Chat [19] 6 4.71 6.06 4.88 5.53 4.78 4.11 4.95 3.03 3.94 2.58 4.88 6.35 6.17 3.94
Yi-VL [3] 9 4.24 4.84 4.74 5.21 3.28 3.57 4.47 2.46 3.27 2.04 4.72 6.62 5.71 4.16
Phi-3-Vision [1] - 3.90 3.99 3.49 4.49 3.87 3.34 3.25 2.32 4.36 2.54 4.05 4.78 5.78 4.42
7 Conclusion
In this paper, we introduce AlignMMBench , a comprehensive Chinese multimodal alignment benchmark comprising three high-level categories and thirteen subtasks. AlignMMBench includes 1,054 images and 4,978 question-answer pairs, encompassing both single-turn and multi-turn dialogue scenarios. To facilitate accurate and efficient evaluations, we developed a critique generation model, referred to as CritiqueVLM . Experimental results demonstrate that CritiqueVLM can assign scores aligned with human preferences and achieve superior performance compared to the widely used GPT-4. Additionally, we present a leaderboard featuring ten popular VLMs supporting Chinese, highlighting potential directions for future improvements in VLMs. We anticipate that this dataset will further advance the development of multimodal language models.
8 Limitations
First, AlignMMBench functions as a benchmark for Chinese language and does not evaluate the multilingual capabilities of VLMs. We plan to gather more images and questions in other languages to extend the scope of our benchmark. Second, due to the suboptimal performance of VLMs in scoring evaluation tasks [7], we currently employ an LLM as our evaluator, which necessitates the use of reference answers that exclude image-derived information. In future work, we will explore the feasibility of using VLMs as reliable evaluators without requiring human-annotated reference answers.
Acknowledgement
We extend our sincere gratitude to the data annotation team at Zhipu AI for their invaluable assistance and support. All GPU compute and API expenses are generously covered by Zhipu AI.
Appendix
Appendix A Datasheet
A.1 Motivation
  • For what purpose was the dataset created? Was there a specific task in mind? Was there a specific gap that needed to be filled? Please provide a description.
    The dataset was created to evaluate the alignment ability of VLMs in Chinese. As far as we know, this is the public first alignment benchmark specifically designed for Chinese VLMs.
  • Who created the dataset (e.g., which team, research group) and on behalf of which entity (e.g., company, institution, organization)?
    The datasets were created by ZhipuAI company.
  • Who funded the creation of the dataset? If there is an associated grant, please provide the name of the grantor and the grant name and number.
    No grant.
  • Any other comments?
    No.
A.2 Composition
  • What do the instances that comprise the dataset represent (e.g., documents, photos, people, countries)? Are there multiple types of instances (e.g., movies, users, and ratings; people and interactions between them; nodes and edges)? Please provide a description.
    The dataset comprises images paired with question-answer pairs. We divide them into 13 different tasks: Description, Recognition, Counting, OCR, Meme, Knowledge, Reasoning, Chart, Problem, Comparison, Writing, Coherence, and Incoherence.
  • How many instances are there in total (of each type, if appropriate)?
    There are 1,054 images and 4,978 question-answer pairs. Detailed distribution in each task type is in our Introduction Section.
  • Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set? If the dataset is a sample, then what is the larger set? Is the sample representative of the larger set (e.g., geographic coverage)? If so, please describe how this representativeness was validated/verified. If it is not representative of the larger set, please describe why not (e.g., to cover a more diverse range of instances, because instances were withheld or unavailable).
    The dataset is a sample from all the Chinese images and user logs.
  • What data does each instance consist of? “Raw” data (e.g., unprocessed text or images) or features? In either case, please provide a description.
    Each instance in the dataset consists of an image, a question, and a reference answer.
  • Is there a label or target associated with each instance? If so, please provide a description.
    Yes, each image-question pair is associated with a reference response that serves as the target for this instance.
  • Is any information missing from individual instances? If so, please provide a description, explaining why this information is missing (e.g., because it was unavailable). This does not include intentionally removed information, but might include, e.g., redacted text.
    No.
  • Are relationships between individual instances made explicit (e.g., users’ movie ratings, social network links)? If so, please describe how these relationships are made explicit.
    N/A.
  • Are there recommended data splits (e.g., training, development/validation, testing)? If so, please provide a description of these splits, explaining the rationale behind them.
    No. The entire AlignMMBench dataset is used to test VLMs’ ability, so there are no recommended data splits.
  • Are there any errors, sources of noise, or redundancies in the dataset? If so, please provide a description.
    Yes, the reference answers are annotated by humans and may contain variations due to the diversity of human expression.
  • Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)? If it links to or relies on external resources, a) are there guarantees that they will exist, and remain constant, over time; b) are there official archival versions of the complete dataset (i.e., including the external resources as they existed at the time the dataset was created); c) are there any restrictions (e.g., licenses, fees) associated with any of the external resources that might apply to a dataset consumer? Please provide descriptions of all external resources and any restrictions associated with them, as well as links or other access points, as appropriate.
    The dataset is self-contained. We gather images from the web and real-user logs, and generate questions and responses through human annotation.
  • Does the dataset contain data that might be considered confidential (e.g., data that is protected by legal privilege or by doctor–patient confidentiality, data that includes the content of individuals’ non-public communications)? If so, please provide a description.
    No.
  • Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety? If so, please describe why.
    No.
  • Any other comments?
    No.
A.3 Collection Process
  • How was the data associated with each instance acquired? Was the data directly observable (e.g., raw text, movie ratings), reported by subjects (e.g., survey responses), or indirectly inferred/derived from other data (e.g., part-of-speech tags, model-based guesses for age or language)? If the data was reported by subjects or indirectly inferred/derived from other data, was the data validated/verified? If so, please describe how.
    The data associated with each instance are annotated by researchers and professional annotators at Zhipu AI.
  • What mechanisms or procedures were used to collect the data (e.g., hardware apparatus(es) or sensor(s), manual human curation, software program(s), software API(s))? How were these mechanisms or procedures validated?
    We employ web crawlers to gather images from the web, utilize the GLM4 API to expand the question set, and let annotators at Zhipu AI annotate reference answers.
  • If the dataset is a sample from a larger set, what was the sampling strategy (e.g., deterministic, probabilistic with specific sampling probabilities)?
    We manually select images to ensure high-quality images paired with question-answer sets.
  • Who was involved in the data collection process (e.g., students, crowdworkers, contractors) and how were they compensated (e.g., how much were crowdworkers paid)?
    Image collection and question construction are carried out by researchers. And Zhipu AI’s professional annotators annotate reference answers. We ensure that annotators are paid more than the minimum wage in our country.
  • Over what timeframe was the data collected? Does this timeframe match the creation timeframe of the data associated with the instances (e.g., recent crawl of old news articles)? If not, please describe the timeframe in which the data associated with the instances was created.
    N/A.
  • Were any ethical review processes conducted (e.g., by an institutional review board)? If so, please provide a description of these review processes, including the outcomes, as well as a link or other access point to any supporting documentation.
    No.
  • Any other comments?
    No.
A.4 Preprocessing/cleaning/labeling
  • Was any preprocessing/cleaning/labeling of the data done (e.g., discretization or bucketing, tokenization, part-of-speech tagging, SIFT feature extraction, removal of instances, processing of missing values)? If so, please provide a description. If not, you may skip the remaining questions in this section.
    Yes. Each image underwent rigorous manual screening to ensure suitability for public use and to avoid inclusion in known SFT datasets or prior benchmarks. Researchers review preliminary annotations, and filter the low-quality annotations.
  • Was the “raw” data saved in addition to the preprocessed/cleaned/labeled data (e.g., to support unanticipated future uses)? If so, please provide a link or other access point to the “raw” data.
    No.
  • Is the software that was used to preprocess/clean/label the data available? If so, please provide a link or other access point.
    No.
  • Any other comments?
    No.
A.5 Uses
  • Has the dataset been used for any tasks already? If so, please provide a description.
    Not beyond this paper.
  • Is there a repository that links to any or all papers or systems that use the dataset? If so, please provide a link or other access point.
    Yes. We will update it on https://alignmmbench.github.io/.
  • What (other) tasks could the dataset be used for?
    The dataset can be applied to other image-related tasks such as visual grounding, image understanding, and visual question answering.
  • Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses? For example, is there anything that a dataset consumer might need to know to avoid uses that could result in unfair treatment of individuals or groups (e.g., stereotyping, quality of service issues) or other risks or harms (e.g., legal risks, financial harms)? If so, please provide a description. Is there anything a dataset consumer could do to mitigate these risks or harms?
    No.
  • Are there tasks for which the dataset should not be used? If so, please provide a description.
    No.
  • Any other comments?
    No.
A.6 Distribution
  • Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created? If so, please provide a description.
    Yes, the dataset will be published later.
  • How will the dataset be distributed (e.g., tarball on website, API, GitHub)? Does the dataset have a digital object identifier (DOI)?
    The dataset will be made available in JSONL format on HuggingFace. As of now, it does not have a DOI.
  • When will the dataset be distributed?
    The dataset will be distributed in July 2024.
  • Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)? If so, please describe this license and/or ToU, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms or ToU, as well as any fees associated with these restrictions.
    Yes, under CC BY-NC-SA 4.0 license.
  • Have any third parties imposed IP-based or other restrictions on the data associated with the instances? If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms, as well as any fees associated with these restrictions.
    No.
  • Do any export controls or other regulatory restrictions apply to the dataset or to individual instances? If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any supporting documentation.
    No.
  • Any other comments?
    No.
A.7 Maintenance
  • Who will be supporting/hosting/maintaining the dataset?
    The dataset will be hosted on HuggingFace.
  • How can the owner/curator/manager of the dataset be contacted (e.g., email address)?
    Email address:
  • Is there an erratum? If so, please provide a link or other access point.
    No.
  • Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)? If so, please describe how often, by whom, and how updates will be communicated to dataset consumers (e.g., mailing list, GitHub)?
    Yes, the dataset will be updated when errors in responses are reported. In the future, we may also add new instances or remove existing ones to ensure the dataset’s accuracy and relevance.
  • If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances (e.g., were the individuals in question told that their data would be retained for a fixed period of time and then deleted)? If so, please describe these limits and explain how they will be enforced.
    N/A.
  • Will older versions of the dataset continue to be supported/hosted/maintained? If so, please describe how. If not, please describe how its obsolescence will be communicated to dataset consumers.
    Yes, we will use HuggingFace to maintain version history.
  • If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so? If so, please provide a description. Will these contributions be validated/verified? If so, please describe how. If not, why not? Is there a process for communicating/distributing these contributions to dataset consumers? If so, please provide a description.
    No.
  • Any other comments?
    No.
Appendix B Radar chart
@@
Appendix C Image examples
In this section, we randomly select multiple images to illustrate the vision diversity of each task, shown as Figure 5.
Figure 5: Image examples of different tasks.
Appendix D Prompt rewrite strategy
@@
Appendix E Evaluation prompt

### v3

@@ A.1 Detailed category definition
  • Incoherence. When a model provides incorrect or irrelevant responses, users usually highlight these errors and expect the model to correct their replies. Therefore, this task assesses the model’s self-correcting abilities.
Model Size Backbone/version Main Language Avg Scores
Qwen2-VL 72B Qwen2-72B+QwenViT Chinese 6.51
CogVLM2 19B EVA2-CLIP-E + Adapter + MLP + LLaMA3 with vision Expert Chinese 5.81
InternVL-V1.5 26B ViT-L/14 + InternLM2 Chinese 5.62
InternVL2 76B InternViT-6B + Llama-3-70B-Instruct Chinese 5.57
MiniCPM-V2.5 8B SigLIP + resampler + LLaMA3 Chinese 5.42
Qwen-VL-Chat 9B Vit-BigG + CrossAttn + Qwen Chinese 5.13
XComposer2V 7B ViT-L/14 + InternLM2-7B Chinese 4.97
DeepSeek-VL 7B SAM-B & SigLIP-L + Adapter + DeepSeek Chinese 4.70
Monkey-Chat 9B Vit-BigG + resampler + Qwen Chinese 4.70
ShareGPT 4V 13B ViT-L/14 + MLP + LLaMA2 English 4.39
LLava1.5 13B ViT-L/14 + MLP + LLaMA2 English 4.31
Yi-VL 34B ViT-H/14 + MLP + Yi Chinese 4.25
Phi3 Vision 4B Vit-L/14 + Phi-3-mini English 4.08
InstructBlip 9B ViT-g/14 + QFormer + LLaMA-7B English 3.31
Table 4: Summary table of different open-source model backbones.
A.2 Radar chart
@@
(a) Distribution of questions.
(b) Distribution of reference answers.
Figure 6: Length distribution of questions and reference answers.
A.3 Prompt rewrite strategy
@@
A.4 Analysis on question and reference answer lengths
In this section, we present the distribution of length for questions and reference answers in our AlignMMBench dataset. Results are illustrated in Figure 6.
A.5 More details of data generation
A.5.1 Image collection
We conducted this process in the following steps: 1) Researchers manually defined 13 task types, and construct keywords for each tasks. 2) Web crawlers collected over 20k images using these keywords and removed duplicates based on their hash values. 3) Researchers manually selected images varying in style and content.
Below are several examples of the keywords in step 1:
  • Description: Poster, Scenery, Flowchart, Map, Cartoon.
  • Knowledge: Movie, Celebrity, Software icon, Famous attraction, Book.
  • Chart: Flowchart, Subway map, Bar graph, Table, Index status, Data summary table.
  • Recognition: Fruit, Flower, Plant, Animal, Food, Simple drawing.
  • OCR: Handwritten, Documents, Signs, Billboards, Math formulas.
  • Counting: Fruit stall, Collections, Crowds.
  • Reasoning: Puzzles, IQ test, Reasoning, Brain teaser.
  • Problem: Math exercises, Reading comprehension problems, Mechanical problems, Biology problems.
  • Comparison: Car comparison, Home products comparison, Product comparison, Cosmetics comparison.
  • Writing: Scenery, Countryside, Forest, Urban, Wildlife.
  • Meme: Meme.
A.5.2 Human annotation
Guidelines
The guidelines provided to human annotators are fundamentally consistent with the prompt detailed in A.7. Furthermore, annotators were supplied with typical characteristics and illustrative examples corresponding to different scoring levels.
Participants and workload
The annotators participating in out work are either employees or interns at Zhipu AI. Prior to initiating the annotation process for this specific project, the annotators participated in an initial trial phase. During this phase, they conducted comprehensive discussions with the project leaders to meticulously refine the annotation guidelines and ensure a comprehensive understanding of the task requirements.
The selection of annotators was not limited to individuals from computer science-related disciplines. Although some researchers and a subset of quality inspectors possess significant expertise in computer science, the general annotators were drawn from a variety of academic and professional backgrounds without stringent restrictions or preferences. This diversity serves to mitigate potential educational bias and aids in creating an alignment benchmark that more accurately represents general human preferences.
Approximately ten individuals are involved in the annotation process, and each annotator processes about 60 entries per day. To ensure the high quality of the annotation results, additional personnel perform quality checks on the annotations at the end of each day.
A.6 Examples
A.6.1 Image examples
In this section, we randomly selecte multiple images to demonstrate the diversity of visual tasks, as depicted in Figure 7. These images illustrate that our AlignMMBench includes a variety of examples that emphasize unique features of the Chinese context. These include: 1) images related to Optical Character Recognition (OCR) and 2) the identification and description of Chinese celebrities, tourist attractions, and holidays. This suggests that our AlignMMBench serves as an excellent supplement to the predominantly English benchmarks currently available.
Figure 7: Image examples of different tasks.
@@ A.6.1 Image examples
A.6.2 Evaluation examples
To more effectively illustrate the evaluation capabilities of CritiqueVLM in comparision to GPT-4, we have chosen an evaluation example in Figure  10.
Figure 10: A specific comparative example between CritiqueVLM and GPT4 evaluation.
A.7 Evaluation prompt